OpenStudy (anonymous):

Bridget has a kitchen canister, in the shape of a cylinder, which holds flour. If the canister is 24 centimeters wide and 26 centimeters tall, what is its approximate surface area in terms of pi? Answer 624π square centimeters 912π square centimeters 1,248π square centimeters 1,296π square centimeters

OpenStudy (anonymous):

get r = 24/2 \[A = 2 \Pi r ^{2} + 2\Pi r h\]

OpenStudy (anonymous):

find (2r^2 + rh)

OpenStudy (anonymous):

i keep getting 2,865 -__________-

OpenStudy (anonymous):

that's because you are multiplying Π=3.14. 2865/Π = 912 which is your answer

OpenStudy (anonymous):

Your supposed to get it in terms of pi

OpenStudy (anonymous):

Ohhhhhh I get it now. Thank you : )

OpenStudy (anonymous):

you're welcome